Thu 8/18/2022 2pm
- Detector status:
- Naomi will email Nick to ask to look at the 2 HVBs that keep tripping
- Beni needs to do a DAQless/direct trigger checkout and then we need to get some cosmics data before the run starts, with time to spare so that we can get problems fixed.
- FA125 - the transient itrig alarm at the start of runs seems to occur less frequently now.
- CDC dE/dx & gain - quantifying the effect of good/bad gain calibrations Naomi made some straightforward estimates of how the count of proton candidates changes as a result of bad gain calibrations. CORRECTION: the gain varies +/- 15% due to changes in atmospheric pressure (not +/- 30%, sorry, bad memory)
- The two GEM trackers for the PS are ready, will discuss with Tim their installation during the downtime. For the GEM-TRD prototype the GEM foil order (from CERN) is on track, some delays with the frames from a Belgium company.